# Codes Correcting Transpositions of Consecutive Symbols

### v1

@@ Abstract
The problem of correcting transpositions (or swaps) of consecutive symbols inqq-ary strings is studied. A family of codes correcting a transposition at an arbitrary location is described and proved to have asymptotically optimal redundancy. Additionally, an improved construction is given over a binary alphabet. Bounds on the cardinality of codes correcting $t=\textrm{const}$ transpositions are obtained. A lower bound on the achievable asymptotic rate of optimal codes correcting $t=\tau n$ transpositions is derived. Finally, a construction of codes correcting all possible patterns of transpositions is presented, and the corresponding lower bound on the zero-error capacity of the $q$-ary transposition channel is stated.
Keywords:
error correction bounds on codes zero-error adjacent transposition spelling error timing error bit shift permutation channel multivariate analytic combinatorics
MSC 2020: 94B25, 94B50, 94B65, 68P30, 68R05
@@ 1 Introduction and preliminaries
Communication channels with various types of reordering and synchronization errors have received increased attention from the research community lately, for the most part due to their applications in modeling modern and somewhat unconventional information transmission and storage systems. In this paper we focus on a model where the transmitted $q$-ary strings are potentially altered by transpositions of consecutive symbols. We first introduce a distance function that is appropriate for this model in the sense that it gives a characterization of the corresponding error correction problem (Section 1.1). We then present two families of codes correcting a single transposition – one over an arbitrary and another over a binary alphabet – and show that they have asymptotically optimal redundancies in the regime of growing block-length (Section 2). We further derive bounds on the cardinality of optimal codes correcting an arbitrary (but fixed) number of transpositions, and conjecture that the upper bound is asymptotically achievable up to a multiplicative constant (Section 3). In the linear regime $t=\tau n$, we derive a Gilbert–Varshamov-like lower bound on the rate of optimal codes correcting a given fraction $\tau$ of transpositions (Section 4). Since the code space under study is nonuniform, unlike in most classical settings, the derivation of such lower bounds is nontrivial and requires estimating the average ball cardinality. The methods of multivariate analytic combinatorics are employed for this purpose. Finally, we present a construction of codes correcting all possible patterns of transpositions, and state the corresponding lower bound on the zero-error capacity of the transposition channel that improves the best known bound for $q\in\{3,4\}$ (Section 5).
@@ 1.1 Model description and basic definitions
For a transmitted string $\boldsymbol{x}=(x_{1},\ldots,x_{n})$, we say that a transposition [^1] Transpositions of consecutive symbols are usually called adjacent transpositions in the literature. We refer to them simply as transpositions. No confusion should arise as this is the only kind of transpositions discussed in the paper. at location $k\in\{1,\ldots,n-1\}$ has occurred in the channel if the string obtained at the channel output is $\boldsymbol{y}=(y_{1},\ldots,y_{n})$, where $y_{k}=x_{k+1}$, $y_{k+1}=x_{k}$, and $y_{i}=x_{i}$ for all $i\notin\{k,k+1\}$, i.e., if the symbols $x_{k}$ and $x_{k+1}$ have been swapped. If the transmitted string is altered by $t\geqslant 1$ transpositions, these transpositions are assumed to be applied simultaneously to disjoint pairs of symbols. In other words, if the transpositions occur at locations $k_{1},\ldots,k_{t}$, then it is understood that $k_{j+1}\geqslant k_{j}+2$ for $j=1,\ldots,t-1$. Consequently, every symbol of $\boldsymbol{x}$ can be displaced by at most one position in the channel. Here is an example of a transmitted string $\boldsymbol{x}\in\mathbb{Z}_{4}^{10}$ and the received string $\boldsymbol{y}$ obtained after $t=3$ transpositions (the pairs of symbols that are being swapped are underlined):
@@
$\hookrightarrow\ 1\ 0\$$1\ 0\ 3\ 0\ 2\ 2\ 1\ 2\ =\ \boldsymbol{y}.$ (1b)
We denote by $B(\boldsymbol{x};t)$ the set of all strings that can be obtained by applying $t$ or fewer transpositions, of the kind just described, to the string $\boldsymbol{x}$.
A nonempty subset of $\mathbb{Z}_{q}^{n}$ is called a code of length $n$, and its elements are called codewords. A code $\mathcal{C}$ is said to be able to correct $t$ transpositions if the following condition holds:
@@
We define the distance between $\boldsymbol{x},\boldsymbol{y}\in\mathbb{Z}_{q}^{n}$ by
$d(\boldsymbol{x},\boldsymbol{y})\coloneqq\min_{t,s\colon B(\boldsymbol{x};t)\cap B(\boldsymbol{y};s)\neq\emptyset}t+s,$ (4)
with the convention $\min\emptyset=\infty$. In words, we look at all the ways a common string $\boldsymbol{z}$ can be produced by applying transpositions to both $\boldsymbol{x}$ and $\boldsymbol{y}$, and among them we choose one requiring the smallest possible cumulative number of transpositions. That number is denoted by $d(\boldsymbol{x},\boldsymbol{y})$. If no common string can be produced by applying transpositions to $\boldsymbol{x}$ and $\boldsymbol{y}$, i.e., if $B(\boldsymbol{x};t)\cap B(\boldsymbol{y};s)=\emptyset$ for all $t,s$, we set $d(\boldsymbol{x},\boldsymbol{y})=\infty$.
@@
We note that $B(\boldsymbol{x};t)\subseteq\bar{B}(\boldsymbol{x};t)$, and that this containment relation is in most cases strict.
Lemma 1
Proof
Suppose a string $\boldsymbol{z}$ can be obtained by applying transpositions at locations $k_{1},\ldots,k_{t}$ to the string $\boldsymbol{x}$, and also by applying transpositions at locations $\ell_{1},\ldots,\ell_{s}$ to the string $\boldsymbol{y}$. If $k_{i}=\ell_{j}$, for some $i\in\{1,\ldots,t\}$, $j\in\{1,\ldots,s\}$, then $(z_{k_{i}},z_{k_{i}+1})=(x_{k_{i}+1},x_{k_{i}})=(y_{k_{i}+1},y_{k_{i}})$, and hence a common string $\boldsymbol{z}^{\prime}$ (obtained by applying a transposition at location $k_{i}$ to $\boldsymbol{z}$) can be obtained by applying t1t-1 transpositions (at locations $k_{1},\ldots,k_{i-1},k_{i+1},\ldots k_{t}$) to $\boldsymbol{x}$ and s1s-1 transpositions (at locations $\ell_{1},\ldots,\ell_{j-1},\ell_{j+1},\ldots\ell_{s}$) to $\boldsymbol{y}$. Therefore, when computing $d(\boldsymbol{x},\boldsymbol{y})$, we may, without loss of generality, assume that the sets of transposition locations $\{k_{1},\ldots,k_{t}\}$ and $\{\ell_{1},\ldots,\ell_{s}\}$ are disjoint. Since $k_{t},\ell_{s}\leqslant n-1$, this implies that $t+s\leqslant n-1$, and hence that $d(\boldsymbol{x},\boldsymbol{y})\leqslant n-1$. That this upper bound is attained is shown by the following examples. For even $n$, take $\boldsymbol{x}$ to be the alternating string $(1,0,1,0,\ldots)$ of length $n$, and $\boldsymbol{y}$ the concatenation of $(0)$, an alternating string $(0,1,0,1,\ldots)$ of length n2n-2, and $(1)$. E.g., for $n=6$, $\boldsymbol{x}=(1,0,1,0,1,0)$ and $\boldsymbol{y}=(0,0,1,0,1,1)$. Then the only common string that $\boldsymbol{x}$ and $\boldsymbol{y}$ can produce is the alternating string $\boldsymbol{z}=(0,1,0,1,\ldots)$, obtained by applying $n/2$ transpositions at locations $1,3,\ldots,n-1$ to $\boldsymbol{x}$, and n/21n/2-1 transpositions at locations $2,4,\ldots,n-2$ to $\boldsymbol{y}$. Similarly, for odd $n$, take $\boldsymbol{x}$ to be the concatenation of the alternating string $(1,0,1,0,\ldots)$ of length n1n-1 and $(0)$, and $\boldsymbol{y}$ the concatenation of $(0)$ and the alternating string $(0,1,0,1,\ldots)$. E.g., for $n=5$, $\boldsymbol{x}=(1,0,1,0,0)$ and $\boldsymbol{y}=(0,0,1,0,1)$. $\blacksquare$
@@
Let $\mathcal{C}\subseteq\mathbb{Z}_{q}^{n}$, $\mathcal{C}\neq\emptyset$. If $d_{\min}(\mathcal{C})>2t$, then $\mathcal{C}$ corrects $t$ transpositions.
@@
By (2), the code $\mathcal{C}$ not being able to correct $t$ transpositions means that there exist distinct codewords $\boldsymbol{x},\boldsymbol{x}^{\prime}\in\mathcal{C}$ such that $B(\boldsymbol{x};t)\cap B(\boldsymbol{x}^{\prime};t)\neq\emptyset$. By (4) and (6), this implies that $d_{\min}(\mathcal{C})\leqslant d(\boldsymbol{x},\boldsymbol{x}^{\prime})\leqslant 2t$. $\blacksquare$
The opposite statement is in general not true. For example, let $\mathcal{C}=\{\boldsymbol{x},\boldsymbol{y}\}\subseteq\mathbb{Z}_{2}^{10}$, where $\boldsymbol{x}=(1,0,1,0,0,0,1,0,1,0)$, $\boldsymbol{y}=(0,0,1,1,0,0,0,0,1,1)$. It is easy to see that the only common string that can be produced by applying transpositions to $\boldsymbol{x}$ and $\boldsymbol{y}$ is $\boldsymbol{z}=(0,1,0,1,0,0,0,1,0,1)$, and that the only way this can be done is by applying transpositions at locations $1,3,7,9$ to $\boldsymbol{x}$, and transpositions at locations $2,8$ to $\boldsymbol{y}$. This implies that $d(\boldsymbol{x},\boldsymbol{y})=4+2=6$, and also that $B(\boldsymbol{x};3)\cap B(\boldsymbol{y};3)=\emptyset$. In other words, the code $\mathcal{C}$ corrects $t=3$ transpositions, and yet $d_{\min}(\mathcal{C})\leqslant 2t$.
@@ 1.2 A related model – successively performed transpositions
We briefly discuss here another model that is closely related to the one introduced in the previous subsection, and for which some of the results of this paper hold by an easy extension.
Namely, if the transmitted string is distorted by $t$ transpositions in the channel, one can imagine a scenario wherein these transpositions are performed in a successive manner, one after another, on pairs of symbols that are not necessarily disjoint. Here is an example with $\boldsymbol{x},\boldsymbol{y}\in\mathbb{Z}_{4}^{10}$ and $t=3$:
@@
In this model, pairs of symbols that are being swapped in two different steps may overlap, and thus one particular symbol of the transmitted string may end up up to $t$ positions away from its original position (see (7a)–(7c)).
@@
The model with successive transpositions is stronger than the one with simultaneous, disjoint transpositions, in the sense that all error patterns that can be realized in the latter can also be realized in the former, i.e., $B(\boldsymbol{x};t)\subseteq B^{\textnormal{{\tiny(s)}}}(\boldsymbol{x};t)$. (In fact, the stronger statement $\bar{B}(\boldsymbol{x};t)\subseteq B^{\textnormal{{\tiny(s)}}}(\boldsymbol{x};t)$ is also true.) Consequently: 1.) any code construction valid for the former model is also valid for the latter, 2.) any lower bound on the cardinality of optimal codes in the former model is also valid in the latter, and 3.) any upper bound on the cardinality of optimal codes in the latter model is also valid in the former. We note that our constructions presented in Section 2 are trivially valid in both models because they concern a single transposition only, in which case there is no difference between the models. Also, the upper bound on the cardinality of optimal codes correcting $t$ disjoint transpositions presented in Theorem 3.1 is, by the above comment 3.), automatically valid for the stronger model with successive transpositions as well.
@@ 1.3 Prior work
Binary codes correcting a fixed number of successive transpositions were analyzed in [16]. Constructions of binary codes correcting a single transposition appear in some earlier works as well, e.g., in [20, 22]. We note that transpositions in binary strings are also known as bit-shifts (or peak-shifts) in the literature and are of relevance in modeling errors in magnetic recording devices. Lower bounds on achievable rates of binary codes correcting a given fraction of bit-shifts were studied in [14, 9].
Zero-error codes, i.e., codes correcting all possible patterns of transpositions, for the model studied herein (as well as for the more general model where each symbol can be displaced by at most $r$ positions) were studied in [21, 2]. Binary zero-error codes for the bit-shift channel (with an additional assumption that each 11 can be shifted by at most $\ell$ positions) were studied in [26, 19]. Zero-error codes for closely related timing channels were studied in [17, 18].
Codes for models that include other types of errors (such as deletions and substitutions) in addition to transpositions were studied in [1, 5, 6, 7, 11, 12, 25, 29, 30].
We emphasize that codes for the $q$-ary transposition channel, which are the focus of the present paper, have not been thoroughly analyzed in the literature. There are only several papers dealing with related problems, e.g., [1, 2, 3, 21, 30].

### v2

# Bounds on Codes Correcting Transpositions of Consecutive Symbols

@@ Abstract
The problem of correcting transpositions (or swaps) of consecutive symbols in $q$-ary strings is studied. Lower bounds on asymptotically achievable rates of codes correcting $t=\tau n$ transpositions are derived. The first bound is obtained by analyzing the average cardinality of “transposition balls” and evaluating the appropriate version of the generalized Gilbert–Varshamov bound, while the second bound follows from a construction of codes correcting an arbitrary number of transpositions (i.e., zero-error codes). Asymptotic bounds on the cardinality of optimal codes correcting $t=\textrm{const}$ transpositions are also derived.
@@ 1 Introduction and preliminaries
Communication channels with various types of reordering and synchronization errors have received increased attention from the research community lately, for the most part due to their applications in modeling modern and somewhat unconventional information transmission and storage systems. In this paper we focus on a model where the transmitted strings are potentially altered by transpositions (or swaps) of consecutive symbols. These and similar kinds of impairments occur in various contexts, e.g., as spelling errors [7], timing and synchronization errors such as packet reordering in communication networks [3] or transpositions in molecular communications [14], intersymbol interference errors such as peak-shifts in magnetic recording channels [29] or charge drift in flash memories [2], genome rearrangement errors [6] which are of relevance for DNA-based data storage systems [9], etc.
In the remainder of this section we describe and discuss the model we are interested in, as well as another closely related model. In particular, we introduce a distance function that is appropriate for our scenario in the sense that it gives a characterization of the corresponding error correction problem (Section 1.1). We then analyze the effects of transpositions on $q$-ary strings and the cardinalities of appropriately defined balls (Section 2). We further derive lower bounds on the largest achievable asymptotic rate of codes correcting $t=\tau n$ transpositions. The first bound is a generalized Gilbert–Varshamov bound (Section 3.1). The second bound is a lower bound on the zero-error capacity of the transposition channel and follows from constructions of codes correcting all possible patterns of transpositions, i.e., zero-error codes. One such construction is presented herein and is shown to improve the best known constructions for $q=3,4$ (Section 3.2). Finally, we obtain asymptotic bounds on the cardinality of optimal codes correcting $t=\textrm{const}$ transpositions (Section 4).
@@ 1.1 Model description and basic definitions
For a transmitted string $\boldsymbol{x}=(x_{1},\ldots,x_{n})$, we say that a transposition [^1] Transpositions of consecutive symbols are usually called adjacent transpositions in the literature. We refer to them simply as transpositions. No confusion should arise as this is the only kind of transpositions discussed in the paper. at location $k\in\{1,\ldots,n-1\}$ has occurred in the channel if the string obtained at the channel output is $\boldsymbol{y}=(y_{1},\ldots,y_{n})$, where $y_{k}=x_{k+1}$, $y_{k+1}=x_{k}$, and $y_{i}=x_{i}$ for all $i\notin\{k,k+1\}$, i.e., if the symbols $x_{k}$ and $x_{k+1}$ have been swapped. If $\boldsymbol{x}$ is altered by $t\geqslant 1$ transpositions, it is assumed that these transpositions are applied simultaneously to $\boldsymbol{x}$, and that therefore they affect disjoint pairs of symbols. In other words, if the transpositions occur at locations $k_{1},\ldots,k_{t}$, it is understood that $k_{j+1}\geqslant k_{j}+2$ for $j=1,\ldots,t-1$. Consequently, every symbol of $\boldsymbol{x}$ can be displaced by at most one position in the channel. Here is an example of a transmitted string $\boldsymbol{x}\in\mathbb{Z}_{4}^{10}$ and the received string $\boldsymbol{y}$ obtained after $t=3$ transpositions (the pairs of symbols that are being swapped are underlined):
@@
$\hookrightarrow 10$ 1 $0302212\ =\ \boldsymbol{y}.$
We denote by $B(\boldsymbol{x};t)$ the set of all strings that can be obtained by applying at most $t$ transpositions, of the kind just described, to the string $\boldsymbol{x}$. It will be helpful to also introduce special notation $B^{\textnormal{e}}(\boldsymbol{x};t)$ for the set of all strings that can be obtained by applying exactly $t$ transpositions to $\boldsymbol{x}$ but cannot be obtained by applying fewer than $t$ transpositions to $\boldsymbol{x}$. The set $B(\boldsymbol{x};t)$ is therefore a disjoint union of the sets $B^{\textnormal{e}}(\boldsymbol{x};t^{\prime})$ over all $t^{\prime}\in\{0,1,\ldots,t\}$,
$B(\boldsymbol{x};t)=\bigsqcup_{t^{\prime}=0}^{t}B^{\textnormal{e}}(\boldsymbol{x};t^{\prime}).$ (2)
Note that $B^{\textnormal{e}}(\boldsymbol{x};t)=\emptyset$ for $n<2t$ because the largest possible number of disjoint transpositions in a string of length $n$ is $\lfloor n/2\rfloor$.
A nonempty subset of $\mathbb{Z}_{q}^{n}$ is called a code of length $n$; its elements are called codewords. A code $\mathcal{C}$ is said to be able to correct $t$ transpositions if the following condition holds:
@@
We say $\mathcal{C}$ is able to correct an arbitrary number of transpositions (or that it is a zero-error code) if the relation (3) holds for every $t\geqslant 0$. Since the largest possible number of disjoint transpositions in a string of length $n$ is $\lfloor n/2\rfloor$, this is equivalent to writing
$\forall\boldsymbol{x},\boldsymbol{x}^{\prime}\in\mathcal{C}\qquad\boldsymbol{x}\neq\boldsymbol{x}^{\prime}\;\Rightarrow\;B(\boldsymbol{x};\lfloor n/2\rfloor)\cap B(\boldsymbol{x}^{\prime};\lfloor n/2\rfloor)=\emptyset.$ (5)
In other words, the statements that a code corrects $\lfloor n/2\rfloor$ transpositions and that it corrects an arbitrary number of transpositions are equivalent.
We define the transposition distance between $\boldsymbol{x},\boldsymbol{y}\in\mathbb{Z}_{q}^{n}$ by
$d(\boldsymbol{x},\boldsymbol{y})\coloneqq\min_{r,s\colon B(\boldsymbol{x};r)\cap B(\boldsymbol{y};s)\neq\emptyset}r+s,$ (6)
with the convention $\min\emptyset=\infty$. In words, we look at all the ways a common string $\boldsymbol{z}$ can be produced by applying transpositions to both $\boldsymbol{x}$ and $\boldsymbol{y}$, and among them we choose one requiring the smallest possible cumulative number of transpositions. That number is denoted by $d(\boldsymbol{x},\boldsymbol{y})$. If no common string can be produced by applying transpositions to $\boldsymbol{x}$ and $\boldsymbol{y}$, i.e., if $B(\boldsymbol{x};r)\cap B(\boldsymbol{y};s)=\emptyset$ for all $r,s$, we set $d(\boldsymbol{x},\boldsymbol{y})=\infty$. Again, since the largest possible number of disjoint transpositions in a string of length $n$ is $\lfloor n/2\rfloor$, we have
$d(\boldsymbol{x},\boldsymbol{y})=\infty\quad\Leftrightarrow\quad B(\boldsymbol{x};\lfloor n/2\rfloor)\cap B(\boldsymbol{y};\lfloor n/2\rfloor)=\emptyset.$ (7)
@@
One can think of $\bar{B}(\boldsymbol{x};t)$ as the set of all strings that can be obtained through a two-step process: first apply $r$ transpositions to $\boldsymbol{x}$, and then apply $s$ transpositions to the resulting string, for any $r,s$ satisfying $r+s\leqslant t$. We note that $B(\boldsymbol{x};t)\subseteq\bar{B}(\boldsymbol{x};t)$, and that this containment relation is in most cases strict.
@@
Let $\mathcal{C}\subseteq\mathbb{Z}_{q}^{n}$, $\mathcal{C}\neq\emptyset$. (a) If $d_{\min}(\mathcal{C})>2t$, then $\mathcal{C}$ corrects $t$ transpositions; (b) $\mathcal{C}$ corrects an arbitrary number of transpositions if and only if $d_{\min}(\mathcal{C})=\infty$.
@@
We first show the contrapositive of (a). If the code $\mathcal{C}$ is not able to correct $t$ transpositions, then, by (3), there exist distinct codewords $\boldsymbol{x},\boldsymbol{x}^{\prime}\in\mathcal{C}$ such that $B(\boldsymbol{x};t)\cap B(\boldsymbol{x}^{\prime};t)\neq\emptyset$. By (6) and (9), this implies that $d_{\min}(\mathcal{C})\leqslant d(\boldsymbol{x},\boldsymbol{x}^{\prime})\leqslant 2t$.
The statement (b) follows from (5) and (7). $\blacksquare$
The opposite statement of (a) is in general not true. For example, let $\mathcal{C}=\{\boldsymbol{x},\boldsymbol{y}\}\subseteq\mathbb{Z}_{2}^{10}$, where $\boldsymbol{x}=(1,0,1,0,0,0,1,0,1,0)$, $\boldsymbol{y}=(0,0,1,1,0,0,0,0,1,1)$. It is easy to see that the only common string that can be produced by applying transpositions to $\boldsymbol{x}$ and $\boldsymbol{y}$ is $\boldsymbol{z}=(0,1,0,1,0,0,0,1,0,1)$, and that the only way this can be done is by applying transpositions at locations $1,3,7,9$ to $\boldsymbol{x}$, and transpositions at locations $2,8$ to $\boldsymbol{y}$. This implies that $d(\boldsymbol{x},\boldsymbol{y})=4+2=6$, and also that $B(\boldsymbol{x};3)\cap B(\boldsymbol{y};3)=\emptyset$. In other words, the code $\mathcal{C}$ corrects $t=3$ transpositions, and yet $d_{\min}(\mathcal{C})\leqslant 2t$.
@@ 1.2 A related model – successively performed transpositions
We briefly discuss here another model that is closely related to the one introduced in the previous subsection.
If the transmitted string is distorted by $t$ transpositions in the channel, one can imagine a scenario wherein these transpositions are performed in a successive manner, one after another, on pairs of symbols that are not necessarily disjoint. Here is an example with $\boldsymbol{x},\boldsymbol{y}\in\mathbb{Z}_{4}^{10}$ and $t=3$:
@@
In this model, pairs of symbols that are being swapped in two different steps may overlap, and thus one particular symbol of the transmitted string may end up up to $t$ positions away from its original position.
@@
The model with successive transpositions is stronger than the one with simultaneous, disjoint transpositions, in the sense that all error patterns that can be realized in the latter can also be realized in the former, i.e., $B(\boldsymbol{x};t)\subseteq B^{\textnormal{{\tiny(s)}}}(\boldsymbol{x};t)$. (In fact, the stronger statement $\bar{B}(\boldsymbol{x};t)\subseteq B^{\textnormal{{\tiny(s)}}}(\boldsymbol{x};t)$ is also true.) Consequently: 1.) any code construction valid for the successive transposition model is also valid for the disjoint transposition model, 2.) any lower bound on the cardinality of optimal codes for the successive transposition model is also valid for the disjoint transposition model, and 3.) any upper bound on the cardinality of optimal codes in the disjoint transposition model is also valid in the successive transposition model. Therefore, the upper bound on the cardinality of optimal codes correcting $t$ disjoint transpositions presented in Theorem 4.1 is automatically valid for the stronger model with successive transpositions as well.
1.3 Transpositions vs. substitutions
For $\boldsymbol{x}\in\mathbb{Z}_{q}^{n}$, let $\boldsymbol{\hat{x}}\in\mathbb{Z}_{q}^{n}$ be the string defined by $\hat{x}_{i}=\sum_{j=1}^{i}x_{j}$, $i=1,\ldots,n$, where addition is performed modulo $q$. Equivalently, $x_{i}=\hat{x}_{i}-\hat{x}_{i-1}$, $i=1,\ldots,n$, where $x_{0}=0$ and subtraction is performed modulo $q$. Then it is easy to see that any $t$ transpositions in $\boldsymbol{x}$ become $t$ substitutions in $\boldsymbol{\hat{x}}$. This is true not only for disjoint but for successive transpositions as well. It follows that, if $\widehat{\mathcal{C}}\subseteq\mathbb{Z}_{q}^{n}$ is a code correcting $t$ substitutions, i.e., a code having minimum Hamming distance larger than $2t$, then $\mathcal{C}=\{\boldsymbol{x}\colon\boldsymbol{\hat{x}}\in\widehat{\mathcal{C}}\}$ is a code correcting $t$ transpositions. Consequently, any lower bound on the cardinality of optimal codes correcting $t$ substitutions is automatically a lower bound on the cardinality of optimal codes correcting $t$ transpositions.
Since the transposition model is in this sense weaker than the substitution model, we expect to be able to achieve higher code rates in the former. This is indeed demonstrated in Section 3.1 and illustrated in Figure 1.
@@ 1.4 Prior work
Asymptotically achievable rates of codes correcting $t=\tau n$ transpositions have been studied previously only in a few special cases. In the binary case ($q=2$), transpositions are also known as bit-shifts or peak-shifts in the literature and are of relevance for modeling errors in magnetic recording devices. Lower bounds on achievable rates of binary codes correcting a fraction of $\tau$ (successively performed) bit-shifts were derived in [19, 12]. Over general alphabets, codes correcting all possible patterns of transpositions, for the model studied herein as well as for the more general model where each symbol can be displaced by up to $\ell$ positions, were studied in [26, 4]. These code constructions imply lower bounds on the zero-error capacity of the transposition channel or, equivalently, on the largest achievable rate of codes correcting a fraction of $\tau=\frac{1}{2}$ transpositions. In the present paper we provide lower bounds on achievable rates valid for any $q\geqslant 2$ and any $\tau\in[0,\frac{1}{2}]$.
We mention in this context also the works studying zero-error codes for models that are not equivalent to the model analyzed here, but are closely related to it, namely the binary bit-shift channel (with an additional assumption that each 11 can be shifted by at most $\ell$ positions) [29, 24] and several types of timing channels [22, 23].
As for codes correcting a fixed number of transpositions, constructions and bounds for the binary case were presented in [21] (for the stronger model with successively performed transpositions/bit-shifts). Constructions of binary codes correcting a single transposition appear in some earlier works as well, e.g., in [25, 27]. In the present paper we obtain asymptotic bounds on the cardinality of optimal $q$-ary codes correcting $t$ transpositions, for any $q\geqslant 2$ and $t\geqslant 1$.
Codes for models that include other types of errors in addition to transpositions, e.g., deletions and/or substitutions, were studied in [1, 5, 7, 9, 10, 16, 17, 28, 32, 33].
Finally, let us mention that transposition errors are also of interest in another scenario where the stored data is represented by a permutation on $\{1,\ldots,n\}$. Namely, transposition-correcting codes in the space of all permutations occur naturally in rank modulation schemes for flash memories; see, e.g., [2].